Harry and Gary are traveling to Nashville
to make their fortunes. Harry leaves on the train at 8:00 A.M.
and Gary travels by car, starting at 9:00 A.M. To complete
the 300-mile trip and arrive at the same time as Harry, Gary
travels 10 miles per hour (mph) faster than the train. At
what time will they both arrive in Nashville?

I am supposed to use the quadratic equation, but this is all I have come up with.
300/8x=300/(9x+10)

Answer by stanbon(75887)

Harry leaves on the train at 8:00 A.M.
and Gary travels by car, starting at 9:00 A.M. To complete
the 300-mile trip and arrive at the same time as Harry, Gary
travels 10 miles per hour (mph) faster than the train. At
what time will they both arrive in Nashville?
--------------------------------------
Gary DATA:
distance = 300 miles ; rate = x mph ; time = d/r = 300/x hrs
-----------------------
Harry DATA:
distance = 300 miles ; rate = (x+10) mph ; time = d/r = 300/(x+10) hrs
----------------------------------------
Equation:
Harry time - Gary time = 1 hr
300/(x+10) - 300/x = 1
300x - 300(x+10) = x(x+10)
-3000 = x^2+10x
x^2 + 10x - 3000 = 0
Factor the Quadratic:
(x+60)(x-50) = 0
Positive solution:
x = 50 mph (Gary's rate)
x+10 = 60 mph (Harry's rate)
----------------------
Cheers,
Stan H.
Reply to stanbon@comcast.net